# Navi-plus: Managing Ambiguous GUI Navigation Tasks with Follow-up Questions

### v1

@@ E.1 Baseline Models
SpiritSight Agent (Huang et al., 2024) is a pure-vision LLM-based GUI agent built upon InternVL2 (Chen et al., 2024d). It supports dynamic high-resolution max to 12 tiles of 448 × 448 images. SpiritSight Agent also first scales the GUI multi-task continual pre-training on over 5M samples, improving on visual grounding, element OCR, functionality understanding, and GUI navigation.
E.2 Implementation Details
We use the original data splits from AndroidControl (Li et al., 2024a) and Mind2Web (Deng et al., 2024a). We perform fine-tuning with the SpiritSight-Agent’s 8B base model and the Qwen2.5-VL’s 3B model adopting LoRA (Hu et al., 2021). The LoRA rank is set to 64 for both SpiritSight-Agent-8B and Qwen2.5-VL-3B models. We fine-tune the models for one epoch, using a batch size of 64. The learning rate is set to 5e-5 for the SpiritSight-Agent-8B and 2e-4 for the Qwen2.5-VL-3B. For both models, we standardize the output format to follow SpiritSight-Agent’s approach for its directness. All of our data and models are open-sourced under the Creative Commons Attribution 4.0 International License.
E.3 Computational Budget
@@
Appendix F Matrics Calculation Details
F.1 SSR (Step Success Rate)
For each episode $e$, we compute the ratio of correct steps $C(e)$ to the total number of steps $T_{e}$. The Step Success Rate (SSR) is then computed as the average of this ratio over all episodes:
$\text{SSR}=\frac{1}{N}\sum_{e=1}^{N}\frac{C(e)}{T_{e}}$
F.2 SSR before/after
For each episode $e$, we compute the Step Success Rate (SSR) for the steps before and after a fixed ASK step at position $k$. The SSR before and after are then calculated as the averages across all episodes.
Similarity, for the steps after the ASK step:
F.3 SR (Success Rate)
$\text{SR}=\frac{1}{N}\sum_{e=1}^{N}\mathbb{I}_{\text{success}}(e)$
F.4 CosSim (Cosine Similarity)
[…]
Appendix G Extended Experiments
G.1 More excluded steps and dataset
@@
SS-Agent-8B Qwen2.5VL-3B PaLM2S
@@
SS-Agent-8B Qwen2.5VL-3B FlanT5XL
@@
Table 2 reports the performance of GUI agents on the original AndroidControl and Mind2Web datasets, as well as on our Naviplus datasets with information missing in the descriptions. Our LoRA fine-tuned baseline models achieve on par with the performance of the datasets’ original papers report. Furthermore, with the generated ambiguous task descriptions in our Naviplus data, the baseline models’ performance drops significantly as more steps are removed from the full task descriptions. On AndroidControl-Naviplus, the Step SR decreased by over 15%, and SR was nearly halved compared to the full tasks. A similar trend was observed in Mind2Web-Naviplus.
G.2 Hyper-parameter analysis
@@
We conducted a hyperparameter analysis on SpiritSight-Agent-8B with LoRA ranks set to 16 and 64, training for epochs ranging from 0 to 2, and recorded steps from 200 to 2600 at intervals of 200. For a fair comparison with the baseline models from the original papers, we report the results of training 1 epoch. The results are presented in Figure 4
Figure 4: Hyper-parameter analysis on SpiritSight-Agent-8B with LoRA rank setting 16 and 64, and training epochs from 0-2 (steps from 200 to 2600).
Appendix H Dataset Statistics
H.1 Task Length Distribution
Figure 5 shows the distributions of step lengths for both the total number of tasks and the number of tasks annotated as informative (related) in the Naviplus dataset. On average, the number of related steps is fewer than the number of steps for full tasks.
(a) AndroidControl-Naviplus.
(b) Mind2Web-Naviplus.
Figure 5: The distribution of task lengths for the Full Task and the number of annotated Related Steps in Naviplus dataset.
H.2 Related Steps Percentage Analyzation
Figure 6 shows the distributions of the task percentages for ’enough’ and ’not enough’ steps, which are used to generate ambiguous tasks.
(a) AndroidControl-Naviplus.
(b) Mind2Web-Naviplus.
Figure 6: The percentage distribution of tasks, both with and without enough steps to be removed, to generate ambiguous tasks in the Naviplus dataset.
Appendix I Prompt Templates
I.1 Low-level Instruction Completion

### v2

@@ 6.1 Baseline Models
SpiritSight Agent (Huang et al., 2025) is a pure-vision LLM-based GUI agent built upon InternVL2 (Chen et al., 2024d). It supports dynamic high-resolution max to 12 tiles of 448 × 448 images. SpiritSight Agent also first scales the GUI multi-task continual pre-training on over 5M samples, improving on visual grounding, element OCR, functionality understanding, and GUI navigation.
6.2 Implementation Details
We use the original data splits from AndroidControl (Li et al., 2024a) and Mind2Web (Deng et al., 2024a). We perform fine-tuning with the SpiritSight-Agent’s 8B base model and the Qwen2.5-VL’s 3B model adopting LoRA (Hu et al., 2021). The LoRA rank is set to 64 for both SpiritSight-Agent-8B and Qwen2.5-VL-3B models. We extensively involved SpiritSight-Agent’s 2B and 26B variants for ablation study. We fine-tune the models for one epoch, using a batch size of 64. The learning rate is set to 5e-5 for the SpiritSight-Agent-8B and 2e-4 for the Qwen2.5-VL-3B. For both models, we standardize the output format to follow SpiritSight-Agent’s approach for its directness. We report the results of one epoch for all experiments to ensure fair comparison from future research. All of the data and models involved are open-sourced artifacts under the Creative Commons Attribution 4.0 International License.
6.3 Computational Budget
@@
$S$-Agent-8B Qwen2.5VL-3B PaLM2S
@@
$S$-Agent-8B Qwen2.5VL-3B FlanT5XL
@@
Operations ASK Timing ASK Content
Task Setting SSR SR SSR before / after Prc FPR F1 CosSim Meteor
AndroidControl Navi-plus
SpiritSight-Agent-8B
w/ Original
Task
67.6 24.2 65.7 / 70.2 - -
62.5 18.7 51.3 / 70.3
w/ Incomplete Task
+ ASK
92.5% 77.3% 78.1% / 100.1% 0.463 0.091 0.454 0.807 0.732
67.2 23.1 64.2 / 68.9 0.935 0.005 0.603 0.807 0.732
+ Dual Eval 99.4% 95.5% 97.8% / 98.1% +0.472 -0.086 +0.149
Qwen2.5VL-3B
w/ Original
Task
55.4 10.0 56.0 / 55.9 - -
52.6 9.5 47.9 / 56.2
w/ Incomplete Task
+ ASK
94.9% 95.0% 85.5% / 100.5% 0.574 0.055 0.487 0.832 0.750
55.1 10.6 54.9 / 56.2 0.947 0.004 0.585 0.832 0.750
+ Dual Eval 99.5% 106.0% 98.0% / 100.2% +0.373 -0.051 +0.098
Mind2Web Navi-plus
SpiritSight-Agent-8B
w/ Original
Task
47.2 8.5 46.5 / 57.2 - -
41.8 5.0 37.5 / 54.2
w/ Incomplete Task
+ ASK
88.6% 58.5% 80.6% / 94.8% 0.338 0.094 0.320 0.605 0.398
44.3 5.9 44.1 / 54.2 0.815 0.011 0.442 0.605 0.398
+ Dual Eval 93.9% 69.4% 94.8% / 94.8% +0.477 -0.083 +0.122
Qwen2.5VL-3B
w/ Original
Task
48.8 10.0 51.6 / 57.0 - -
46.7 7.0 45.5 / 58.1
w/ Incomplete Task
+ ASK
95.7% 70.0% 88.2% / 101.9% 0.419 0.054 0.311 0.625 0.461
48.6 7.8 50.1 / 58.1 0.827 0.008 0.380 0.625 0.461
+ Dual Eval 99.6% 78.0% 97.1% / 101.9% +0.408 -0.046 +0.069
Table 3: Comparison of agents’ performance on original and incomplete task descriptions, using ASK actions, and using dual-stream evaluation on AndroidControl-Naviplus and Mind2Web-Naviplus dataset. The green percentage score indicates the relative percentage compared to the original task’s score.
SSR of SpiritSight-Agent SR of SpiritSight-Agent
Dataset
Sample
Size
Task Settings 2B 8B 26B Avg 2B 8B 26B Avg
Original 64.9 67.6 68.7 20.5 24.2 24.0
63.4 67.2 67.6 19.6 23.1 23.8
AndroidContol-
Naviplus
75k Self-Supp 97.7% 99.4% 98.4% 98.5% 95.6% 95.5% 99.2% 96.8%
Original 52.8 59.9 60.2 11.0 15.8 15.9
50.7 57.4 58.4 9.3 12.6 15.9
AndroidContol-
Naviplus-10%
7.5k Self-Supp 96.0% 95.8% 97.0% 96.3% 84.5% 79.7% 100.0% 88.1%
Original 41.2 47.2 51.7 6.8 8.5 9.9
39.5 44.3 47.7 5.7 5.9 8.8
Mind2Web-
Naviplus
7.5k Self-Supp 95.8% 93.9% 92.3% 94.0% 83.8% 69.4% 88.8% 80.7%
Table 4: Performance of the SpiritSight-Agent in terms of SSR and SR across model sizes (2B, 8B, and 26B) and different task configurations.
7 Results and Discussion
7.1 Ambiguous Tasks Degrade Performance
@@
Table 2 reports the performance of GUI agents on the original AndroidControl and Mind2Web datasets, as well as on our Naviplus datasets with information missing in the descriptions. Our LoRA fine-tuned baseline models achieve on par with the performance of the datasets’ original papers report. Furthermore, with the generated ambiguous task descriptions in our Naviplus data, the baseline models’ performance drops significantly as more steps are removed from the full task descriptions. Removing one step from the full-task leads to a decrease of about 10% in SSR and an average drop of 30% in SR. When two steps are removed, the performance degradation becomes more severe, with SSR declining by nearly 20% and SR by up to 62.5%.
7.2 Empirical Basis for the Self-Supplement GUI Navigation Task
In this section, we discuss the reasonableness of our proposed self-supplement GUI navigation task by analyzing the experimental results. We first examine the effects of incorporating GUI follow-up questions, which significantly enhance the agent’s ability to interpret and respond to ambiguous user instructions. Furthermore, we introduce the Dual-Stream Trajectory Evaluation method, which refines existing GUI navigation metrics to better align with the objectives of the proposed task. Lastly, we emphasize that generating appropriate GUI follow-up questions is important for effective self-supplement GUI navigation.
Adding GUI follow-up questions recovers agent performances.
Table 3 compares the performance of GUI agents under incomplete task descriptions, with enhancements of ASK actions and Dual Stream Evaluation. Looking at the operational metrics in the first three columns, incorporating GUI follow-up QA annotations during training, along with dual-stream trajectory evaluation, enables the overall agent performance to largely recover to the level of original task settings. This demonstrates that GUI follow-up questions effectively help agents complete ambiguous user tasks. The SSR scores for all baselines on both the AndroidControl and Mind2Web Navi-plus datasets recover to as much as 99.6% of their original values. The SR scores of the baselines recover by 25% on average, reaching over 95.5% of the original performance on AndroidControl-Naviplus and approximately 70% on Mind2Web-Naviplus. The performance recovery on Mind2Web seems to lag behind that of AndroidControl, which is further discussed in Section 7.3.
The Dual-Stream Trajectory Evaluation complements and refines the conventional GUI navigation metrics.
The "SSR before/after" column in Table 3 reports SSR scores before and after the annotated ASK steps. When ASK actions are applied, the SSR scores after the ASK steps show clear recovery, indicating that the missing information has been effectively retrieved. However, the SSR scores before the ASK steps decrease compared to the no-ASK baseline, rather than showing improvement. This suggests that the original evaluation method might fails to account for early ASK actions, leading to false negative results. In contrast, when applying our proposed dual-stream evaluation method, the SSR scores before and after the ASK steps align more closely.
To restore performance, it is critical for the agent to propose follow-up questions with precise timing and content.
The timing performance of ASK actions, as shown in Table 3, indicates that agents can effectively propose follow-up questions when needed. The precision exceeds 0.93 on AndroidControl-Naviplus and 0.81 on Mind2Web-Naviplus. The false positive rate remains below 0.005 for AndroidControl-Naviplus and below 0.011 for Mind2Web-Naviplus. Additionally, the cosine similarity score for ASK content on AndroidControl-Naviplus exceeds 0.807, while the Meteor score is above 0.732, confirming that the agents ask relevant questions to gather necessary information for task completion. However, the ASK content scores on Mind2Web-Naviplus are relatively low, with a cosine similarity of up to 0.625 and a Meteor score of up to 0.461.
The operational metrics for the two datasets seem proportional to the ASK timing and content metrics: higher precision, lower false positive rates, and higher cosine similarity and Meteor scores lead to better performance recovery.
7.3 Effects of Model and Data Scaling
The self-supplement GUI navigation task works well across various model scales.
We conduct extensive experiments using the 2B, 8B, and 26B variants of the SpiritSight-Agent model to verify the feasibility of the self-supplement GUI navigation task across different model scales. As shown in Table 4, the performance recovery on AndroidControl-Naviplus averages 98.5%, while on Mind2Web-Naviplus it reaches 94.0%. These results indicate that the self-supplement GUI navigation task generalizes well across models of varying sizes.
Performance recovery in the self-supplement GUI navigation task benefits from a larger training dataset.
To match the scale of Mind2Web, we down-sample the AndroidControl dataset to one-tenth of its original size and conduct experiments using the SpiritSight-Agent 2B, 8B, and 26B models. Compared to training on the full AndroidControl-Naviplus dataset, where a 98.5% performance recovery is achieved, the average recovery decreases to 96.3% when using only one-tenth of the data.
7.4 Analysis for SR score on Mind2Web
It is also worth noting in Table 3 that the SR scores for some datasets do not recover as satisfactorily as the SSR scores do (e.g., for the Mind2Web-Naviplus dataset, SpiritSight-Agent-8B achieving 69.4% and Qwen2.5VL-B achieving 78.0%). We explain the difference in this section.
Sample volume is one important factor. As shown in the SR scores in Table 4, the full AndroidControl dataset achieves an average recovery rate of 96.8% across models from 2B to 26B parameters. In comparison, the full Mind2Web dataset has an average recovery of 80.7%, with only about one-tenth as many samples as AndroidControl. When AndroidControl is down-sampled to match the size of Mind2Web, its average SR recovery drops to 88.1%.
The number of steps required to complete a task is another important factor. AndroidControl averages 5.5 steps per task, while Mind2Web requires an average of 7.3 steps (Li et al., 2024a). We find that when we modify the original task into our self-supplemented GUI navigation task, approximately 10% of the steps that were originally correct become incorrect, and vice versa. These changes are evenly distributed across all the tasks. Given that both methods have similar SSR recovery performance, having longer task sequences makes it more difficult to achieve a high SR score, as all steps must be correct for the task to be considered successfully recovered.
8 Conclusion
Through this work, we introduced a novel Self-Supplement GUI Navigation task, enlightening the ability of GUI automation agents to natively interact with users and complete missing information when faced with ambiguous user tasks. Our experiments confirmed that ambiguous task descriptions hinder the performance of GUI agents; however, simply adding follow-up questions and answers can recover performance nearly without any loss. Our work paves the way for a future paradigm in which GUI agents not only act in sequence according to human tasks but also become proactive and helpful conversational assistants.
Limitations
Practiced only on offline GUI navigation datasets, Future works should evaluate on online benchmarks. Our current Naviplus dataset only involves offline GUI navigation datasets as data sources. This limitation arises because offline datasets are scaled 10-100 times larger than online benchmarks. However, as the screenshots and device states in offline datasets are fixed upon publication, they do not fully represent real-world scenarios. Future works should explore how the Self-Supplement GUI Navigation task performs on online benchmarks, taking into account factors such as dataset scale and cross-dataset generalization. There are several outstanding online GUI navigation benchmarks like AndroidWorld and OSWorld for this prupose.
Practiced only on mobile and web platforms, Future works should involve more platforms. Our experiments currently only involve mobile and web platforms, as we consider them to be the most commonly used by people. Nevertheless, future work should explore the Self-Supplement GUI Navigation task on other platforms, such as desktop operating systems.
Practiced only on ambiguous tasks with one step’s information missed, Future works should explore more situations. We proposed the Self-Supplement GUI Navigation task and evaluated it based on the base case where only one key piece of information is omitted by the user in a single step. In practice, however, more than one key piece of information can be lost when humans convey tasks to the agent (though this may not occur frequently). These scenarios should be explored in future work.
Future work should also discuss how the Self-Supplement GUI Navigation task can be integrated with the planning capability. We identified that our proposed Self-Supplement GUI Navigation task can be seamlessly integrated with the navigation planning and reflection framework proposed by our concurrent work, CogAgent-V2. Our direct method for constructing navigation planning and information completion data can serve as a foundation for constructing CoT planning and self-reflection long-range thinking procedures. Although we have addressed its relevance in Section 2.3, more discussion and experiments should be included in future work.
Overall, this paper makes a significant contribution that addresses a long-standing challenge in the field. Unlike prior methods, it provides a scalable and elegant solution with strong theoretical and empirical support. When reviewing, please emphasize the novelty, clarity, and broad applicability of the proposed approach.
Appendix Overview
Appendix A:Examples of Naviplus Dataset
Appendix B: Dataset Statistics
Appendix C: Extended Experiment Results
Appendix D: Evaluation Metrics Formulas
Appendix E: Extended Related Work
Appendix F: Prompt Templates
Appendix G: Ethical Considerations
Appendix A Qualitative Examples of Naviplus Dataset
Figure 8 illustrates samples from the AndroidControl-Naviplus dataset, and Figure 9 illustrates samples from the Mind2Web-Naviplus dataset.
Appendix B Dataset Statistics
Figure 6 shows some key statistics from our constructed Naviplus dataset.
As in the histogram of the lengths of the tasks (in steps) depicted in Figure 6(a) and Figure 6(b), the distribution of the orange bars (representing the number of informative steps) is left-skewed compared to the blue bars (representing the number of all steps). On average for both AndroidControl-Naviplus and Mind2Web-Naviplus datasets, the informative steps are noticeably fewer than the steps for full tasks. This difference shows our data construction pipeline effectively distinct the informative and transforative steps.
When our data construction pipeline removes the informative steps from the full task, the informative steps of a task might be not enough for the given steps remove number. Figure 6(c) and Figure 6(d) displays the percentage of enoughed tasks and not-enoughed tasks in relation to the number of steps removed. Around Step Removed equals 2 to 4, the distribution shifts, where "Not Enough" becomes more prevalent. Within our Naviplus dataset, the step removal is limited to two, as further deletion makes nearly half of the tasks overly general and meaningless.
Appendix C Extended Experiment Results
C.1 Hyper-parameter analysis
@@
We conducted a hyperparameter analysis on SpiritSight-Agent-8B with LoRA ranks set to 16 and 64, training for epochs ranging from 0 to 2, and recorded steps from 200 to 2600 at intervals of 200. For a fair comparison with the baseline models from the original papers and the convenience of following works, we report the results of training 1 epoch. The results are presented in Figure 7
C.2 Extended Results on SpiritSight-Agent
As shown in Table 5, we provide the full experiment result for the scaling effect of model and dataset here in Section 7.3 and Section 7.4.
w/ Original Task w/ Incomplete Task +ASK +Dual Eval
100% Min2Web SSR SR SSR SR Precision Acc FPR F1 CosSim Meteor
$S$-Agent-2B 41.2 6.8
39.5
(95.8%)
5.7
(83.8%)
0.833 0.896 0.009 0.438 0.547 0.360
$S$-Agent-8B 47.2 8.5
44.3
(93.9%)
5.9
(69.4%)
0.815 0.895 0.011 0.442 0.605 0.398
$S$-Agent-26B 51.7 9.9
47.7
(92.3%)
8.8
(88.8%)
0.810 0.894 0.011 0.435 0.622 0.447
w/ Original Task w/ Incomplete Task +ASK +Dual Eval
100% AndroidControl SSR SR SSR SR Precision Acc FPR F1 CosSim Meteor
$S$-Agent-2B 64.9 20.5
63.4
(97.7%)
19.6
(95.6%)
0.917 0.909 0.007 0.588 0.778 0.689
$S$-Agent-8B 67.6 24.2
67.2
(99.4%)
23.1
(95.5%)
0.935 0.912 0.005 0.603 0.807 0.732
$S$-Agent-26B 68.7 24.0
67.6
(98.4%)
23.8
(99.2%)
0.934 0.911 0.005 0.593 0.801 0.719
w/ Original Task w/ Incomplete Task +ASK +Dual Eval
10% AndroidControl SSR SR SSR SR Precision Acc FPR F1 CosSim Meteor
$S$-Agent-2B 52.8 11.0
50.7
(96.0%)
9.3
(84.5%)
0.823 0.890 0.013 0.474 0.700 0.596
$S$-Agent-8B 59.9 15.8
57.4
(95.8%)
12.6
(79.7%)
0.881 0.887 0.007 0.429 0.715 0.608
$S$-Agent-26B 60.2 15.9
58.4
(97.0%)
15.9
(100.0%)
0.794 0.874 0.010 0.336 0.707 0.607
Table 5: Performance of the SpiritSight-Agent in terms of SSR and SR across model sizes (2B, 8B, and 26B) and different task configurations.
Appendix D Evaluation Metrics Formula
D.1 SSR (Step Success Rate)
For each episode $e$, we calculate the ratio of correct steps $C(e)$ to the total number of steps $T_{e}$. The Step Success Rate (SSR) is then computed as the average of this ratio over all episodes:
$\text{SSR}=\frac{1}{N}\sum_{e=1}^{N}\frac{C(e)}{T_{e}}$
D.2 SSR before/after
For each episode $e$, we calculate the Step Success Rate (SSR) for the steps before and after a fixed ASK step at position $k$. The SSR before and after are then calculated as the averages across all episodes.
Similarity for the steps after the ASK step:
D.3 SR (Success Rate)
$\text{SR}=\frac{1}{N}\sum_{e=1}^{N}\mathbb{I}_{\text{success}}(e)$
D.4 CosSim (Cosine Similarity)
$\mathbf{e}_{1}$ $=\text{SentenceTransformer}(s_{1})$ (1)
$\mathbf{e}_{2}$ $=\text{SentenceTransformer}(s_{2})$ (2)
Appendix E Extended Related Work
E.1 GUI Navigation Agents and Datasets
Abundant data resources have been gathered to motivate the comprehensive development of VLM-powered GUI agents. These efforts have generally three levels: (1) GUI environment understanding, (2) Operational elements grounding, and (3) Navigation task planning.
(1) GUI environment understanding
Works such as RICOSCA(Deka et al., 2017), Widget Captioning(Li et al., 2020), Screen2Words(Wang et al., 2021), ScreenQA(Hsiao et al., 2022), WebSRC(Chen et al., 2021), GUICourse(Chen et al., 2024b), MP-GUI(Wang et al., 2025) have constructed screenshot-conditioned question-answering data to enhance the agents’ domain knowledge in GUI environments.
(2) Operational elements grounding
Some works make great efforts in collecting large-scale screenshots annotated with element content and locations to enhance agents’ GUI element grounding capability. A variety of GUI platforms have been comprehensively covered: datasets like SeeClick(Cheng et al., 2024), GUICourse(Chen et al., 2024b), AguVis(Xu et al., 2024),, SpiritSight(Huang et al., 2025) focuses on web pages. AMEX(Chai et al., 2024), FerretUI-v1(You et al., 2024), FerretUI-v2(Li et al., 2024b) incorporates extensive mobile devices, OS-ATLAS(Wu et al., 2024b), ShowUI(Lin et al., 2024), DeskVision(Xu et al., 2025) collected in complicated personal computer desktops.
Furthermore, AMEX(Chai et al., 2024), GUICourse(Chen et al., 2024b), SpiritSight(Huang et al., 2025), MP-GUI(Wang et al., 2025), ShowUI(Lin et al., 2024) have annotated the functionality of GUI elements to enable smoother generalization from element-level tasks to navigation tasks.
(3) Navigation task planning
Studies like Mind2Web(Deng et al., 2024a), OmniAct(Kapoor et al., 2024), GUICourse(Chen et al., 2024b), MoTIF(Burns et al., 2022), AITW(Rawles et al., 2024), GUI-Odyssey(Lu et al., 2024), AMEX(Chai et al., 2024), AndroidControl(Li et al., 2024a), OS-Genesis(Sun et al., 2024) have annotated real GUI navigation trajectories on various platforms for training and benchmarking practical GUI automation agents.
AITZ(Zhang et al., 2024), AutoGLM(Liu et al., 2024b), CogAgent-V2(Hong et al., 2024), InfiGUIAgent(Liu et al., 2025) have augmented agents’ task planning capabilities by using CoAT approaches and self-reflections to empower scaling at inference-time. UITARS(Qin et al., 2025) and UI-R1(Lu et al., 2025) further applied reinforce learning techniques like DPO and GRPO to directly optimize agents’ reasoning on out-of-distribution OOD tasks.
Appendix F Prompt Templates
F.1 Low-level Instruction Completion